OpenStudy (anonymous):

HELP: write in slope intercept form: Slope = 1/2 through (-2,8) I know this far; y= 1/2x ...

OpenStudy (anonymous):

y = mx + c, where m is the slope or gradient and c is the intercept point. use slope = 1/2 to give y = 1/2 x + c and then use the points coordinates to find what c is